OpenStudy (anonymous):

Solve triangle UVW if m angle U = 41°, w = 12, and v = 13. A) m angle U = 41°, m angle V ≈ 75.7°, m angle W≈ 63.3°, and u ≈ 8.8, v = 13, and w = 12 B) m angle U = 41°, m angle V ≈ 63.3°, m angle W ≈ 75.7°, and u ≈ 9.9, v = 13, and w = 12 C) m angle U = 41°, m angle V ≈ 63.3°, m angle W ≈ 75.7°, and u ≈ 8.8, v = 13, and w = 12 D) m angle U = 75.7°, m angle V ≈ 41°, m angle W ≈ 63.3°, and u ≈ 8.8, v = 12, and w = 13

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

wow this involves a lot of work lol hold on :P

OpenStudy (anonymous):

sorry lol i just dont understand it

HanAkoSolo (jamierox4ev3r):

its basically the same process as the other problems we've done. It just involves solving for more things T_T

OpenStudy (anonymous):

yea.. :/

HanAkoSolo (jamierox4ev3r):

I'm just going to change uvw to abc if you don't mind...makes things easier for me

OpenStudy (anonymous):

yea thats fine

HanAkoSolo (jamierox4ev3r):

alright wow this problem though...tedious as heck! btw, we can eliminate option D right off the bat, it contradicts the first part of the problem lol

OpenStudy (anonymous):

\[u=\sqrt{w^2+v^2-2wvcos(U)}\]

HanAkoSolo (jamierox4ev3r):

you can solve for side a (or u) first

OpenStudy (anonymous):

12^2 +13^2 - 2(12)(13)cos(U)

HanAkoSolo (jamierox4ev3r):

yep! and once you solve the side, you can possibly switch to law of sines if you so choose to find the rest!

HanAkoSolo (jamierox4ev3r):

I actually have to go really soon, its really late for me ;-;

OpenStudy (anonymous):

ok just 1 more min plz

OpenStudy (anonymous):

144 + 169 - 312 = 6

OpenStudy (anonymous):

and...she leaves me :/

OpenStudy (anonymous):

i got no answer :c

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Jhannybean

TheSmartOne (thesmartone):

I haven't learned this yet :/ Maybe @Jhannybean @Michele_Laino @perl @Miracrown can help you :)

OpenStudy (michele_laino):

Please I apply the Carnot Theorem, so I get: \[{u^2} = {v^2} + {w^2} - 2vw\cos 41 = 169 + 144 - 312\cos 41\] what is u?

OpenStudy (anonymous):

angle U = 41 degrees

OpenStudy (michele_laino):

yes! 

OpenStudy (anonymous):

ok so whats next

OpenStudy (michele_laino):

next please you have to perform the computation, namely: \[{u^2} = 169 + 144 - 312\cos 41 = 313 - 312 \cdot 0.755 = 313 - 235.6 = 77.4\]

OpenStudy (michele_laino):

so we have: \[{u^2} = 77.4\]

OpenStudy (anonymous):

ok but in the choices its 75.7

OpenStudy (michele_laino):

please you have to look at the values for u

OpenStudy (michele_laino):

so you have to compute this: \[u = \sqrt {77.4} = ...?\]

OpenStudy (anonymous):

ohh so it would be 8.7

OpenStudy (michele_laino):

I think 8.8 since you have to round off your result

OpenStudy (anonymous):

yea true

OpenStudy (anonymous):

ok so now we should find v

OpenStudy (michele_laino):

please note that in order to find V, you can apply the sines Theorem

OpenStudy (michele_laino):

namely: \[\frac{v}{{\sin V}} = \frac{u}{{\sin 41}}\]

OpenStudy (michele_laino):

and substituting the values for u and v, we get: \[\frac{{13}}{{\sin V}} = \frac{{8.8}}{{\sin 41}}\]

OpenStudy (michele_laino):

next you have to solve that equation for sinV, so I can re-write it as follows: \[\sin V = \frac{{13}}{{8.8}}\sin 41\]

OpenStudy (michele_laino):

please, what is sinV?

OpenStudy (anonymous):

would it be 1.32

OpenStudy (michele_laino):

no, I don't think since sinV has to be a number less than 1

OpenStudy (michele_laino):

hint: \[inV = \frac{{13}}{{8.8}}\sin 41 = \frac{{13}}{{8.8}} \cdot 0.656 = ...?\]

OpenStudy (michele_laino):

oops... \[\sin V = \frac{{13}}{{8.8}}\sin 41 = \frac{{13}}{{8.8}} \cdot 0.656 = ...?\]

OpenStudy (anonymous):

0.96

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so what is V?

OpenStudy (michele_laino):

please you have to use your scientific calculator

OpenStudy (anonymous):

i dont know how to get v

OpenStudy (michele_laino):

please you have to use the inverse function option, nevertheless here i your value for V: V=73.74 degrees, namely: V = 74 degrees

OpenStudy (michele_laino):

oops... here is your value...

OpenStudy (anonymous):

oh ok so i think the answer would be A

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

ok thanks for your help :)

OpenStudy (michele_laino):

thank you! :)